Three blocks on a frictionless surface are connected by massless strings, with M1 = 1.70kg, M2 = 2.10kg, and M3 = 3.20kg. Due to the force F acting on M3, as shown, the system accelerates to the right. Given that T1 is 3.50 Newtons, calculate T2.
Three blocks on a frictionless surface are connect
Answer #1

Accelaration = (f/m), = 3.5/1.7, = 2.05m/sec^2. All masses must have the same acceleration.

Force = (ma), so T2 = (1.7 + 2.1) x 2.05, = 7.82N.